OpenStudy (anonymous):

what is the integral from 0 to pi/2 of cos(x)* e^sinxdx

OpenStudy (anonymous):

You should look at this as e^u du..

OpenStudy (campbell_st):

use the substitution u = sin(x) and then du/dx = cos(x) or du = cos(x) dx so yyou have the integral \[\int\limits e^u du\] just do the boundary values

OpenStudy (anonymous):

thank you! @campbell_st